OpenStudy (anonymous):

Can someone show me how to do this on a calculator? 2^x = 14

OpenStudy (rsadhvika):

take log both sides

OpenStudy (rsadhvika):

then you can use log button in calculator to find x

OpenStudy (cwrw238):

yes thats what i have to do on my calculator

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

no need for a calculator when you are on a computer http://www.wolframalpha.com/input/?i=ln%2814%29%2Fln%282%29

OpenStudy (anonymous):

Ok, thank you!